OpenStudy (anonymous):

To the nearest hundredth, what is the radius of a circle with a circumference of 9?

OpenStudy (anonymous):

2*pi*r = 9 r = 9/(2*pi) r = 1.432394488 radius = 1.43 units

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :)